# OccFiner: Offboard Occupancy Refinement with Hybrid Propagation

### v1

@@ Abstract
Vision-based occupancy prediction, also known as 3D Semantic Scene Completion (SSC), presents a significant challenge in computer vision. Previous methods, confined to onboard processing, struggle with simultaneous geometric and semantic estimation, continuity across varying viewpoints, and single-view occlusion. Our paper introduces OccFiner, a novel offboard framework designed to enhance the accuracy of vision-based occupancy predictions. OccFiner operates in two hybrid phases: 1) a multi-to-multi local propagation network that implicitly aligns and processes multiple local frames for correcting onboard model errors and consistently enhancing occupancy accuracy across all distances. 2) the region-centric global propagation, focuses on refining labels using explicit multi-view geometry and integrating sensor bias, especially to increase the accuracy of distant occupied voxels. Extensive experiments demonstrate that OccFiner improves both geometric and semantic accuracy across various types of coarse occupancy, setting a new state-of-the-art performance on the SemanticKITTI dataset. Notably, OccFiner elevates vision-based SSC models to a level even surpassing that of LiDAR-based onboard SSC models.
Keywords:
Semantic Scene Completion 3D Occupancy Prediction Offboard Perception
1 Introduction
@@
Vision-based Occupancy Prediction, often referred to as Semantic Scene Completion (SSC), is a process aimed at accurately reconstructing the geometry and semantics of a surrounding 3D scene from the image captures [6]. Compared to LiDAR solutions, camera-based systems offer the benefits of being lightweight, cost-effective, and ease of deployment and maintenance. Most SSC efforts are made in onboard settings [52, 6, 27], i.e., data calculation and feedback take place on the mounted vehicle. However, integrating data into offboard systems creates opportunities for identifying and rectifying potential deficiencies in models. Consequently, we propose offboard SSC, a new setup that entails additional computation and refinement of on-vehicle predictions at a data center. It can play a pivotal role in auto-labeling by leveraging aggregated and diverse data from multiple sources, significantly reducing manual annotation efforts and improving labeling accuracy for large-scale 3D scene understanding [77, 25, 24].
@@
Despite various research efforts [6, 27, 18], vision-based SSC has fallen short in accuracy compared to LiDAR solutions. There are two limitations: 1) The depth measurement of monocular cameras lacks the precision of LiDAR systems, requiring onboard vision-based SSC methods to struggle with both geometry and semantics. 2) Onboard models are limited to processing a restricted number of local frames, impacting accuracy and multi-view continuity in both vision- and LiDAR-based SSC. This issue manifests as unreliable estimations and discontinuities across viewpoints in geometry and semantics, as illustrated in Fig. 1. We argue that high-fidelity SSC maps should remain consistent across viewpoints to ensure safety and reliability, especially since static rigid objects should not exhibit significant changes in classes. While offboard systems offer refined evaluation of onboard model performance, essential for the data closed-loop in autonomous driving [77, 25, 24], their potential in SSC remains under-explored.
@@
Given such limitations of inferior performance and onboard SSC setting, we introduce OccFiner, the first offboard occupancy refinement framework. The objective is to significantly enhance the reliability of vision-based SSC. By doing so, the constructed 3D SSC map can be reliably used repeatedly after a one-time low-cost acquisition with only cameras. Specifically, we consider two primary error sources in addressing the inaccuracies of vision-based SSC: the prediction bias error introduced by the onboard model, and the measurement bias from the camera capture process. OccFiner decouples these two heterogeneous error sources and uses a two-stage hybrid process to process them separately: (1) Multi-to-Multi Local Propagation. In the initial stage, OccFiner addresses prediction bias errors from the onboard model. We implement an error compensation model that processes inputs from off-the-shelf SSC models [6, 27, 52]. This model uniquely encodes relative spatial coordinates within a local window, facilitating the propagation of geometric and semantic cues by a newly proposed DualFlow4D transformer. This approach effectively enhances the accuracy of occupancy predictions at varying distances. (2) Region-centric Global Propagation. The second stage focuses on measurement errors from camera captures. Here, OccFiner employs explicit multi-view geometry to register and aggregate data across the entire scene, incorporating senor bias weighting in the process. Compared with the simple accumulation of multi-frame averages, our region-centric voting strategy yields a more accurate SSC map. By integrating local implicit error compensation and non-local explicit information propagation from various viewpoints, OccFiner effectively addresses onboard bias, multi-view consistency, and single-view occlusions, therefore naturally enhancing the reliability and robustness of scene understanding.
@@
To the best of our knowledge, we are the first to introduce offboard refinement for SSC. Our method, termed OccFiner, enables seamless integration across various SSC models and datasets as a plug-and-play solution. For 3D semantic scene completion on SemanticKITTI dataset [2], OccFiner has demonstrated remarkable performance, surpassing VoxFormer by $3.67\%$ in mIoU, which makes a $26.99\%$ relative improvement and ranks first on the test leaderboard among all monocular methods. Surprisingly, OccRefiner also works well for LiDAR-based SSC [65] with minor changes, leading to a new state-of-the-art of $37.82\%$ in mIoU, achieving a $9.5\%$ improvement compared with the best benchmark result.
@@
  • We are the first to explore the problem of learning to generate high-quality vision-oriented SSC maps offboard.
  • We propose OccFiner, an effective offboard occupancy refinement framework with hybrid propagation to build multi-view consistent SSC maps.
  • OccFiner sets new state-of-the-art performances in both camera-based and LiDAR-based SSC. For the first time, the accuracy of the camera-based SSC method surpasses that of the classical onboard LiDAR-based method [55].
2 Related Work
2.1 3D Semantic Scene Completion
@@ 2 Related Work
2.2 Offboard 3D Perception
Perception in autonomous driving scenarios necessitates extensive data annotation and training [11, 5, 56]. Since LiDAR can provide accurate range information in 3D space, researchers explore its sequence learning in semantic segmentation [61, 1], object detection [44, 47] and tracking [46] to obtain better performance. The sequential characteristics of LiDAR scans are further studied in offboard 3D object detection [50, 71, 10, 37] to reduce annotation costs. 3DAL [50] formulates the problem of offboard 3D object detection and proposes an object-centric auto-labeling detector. CTRL [10] adheres to the principle of track-centric and elevates the accuracy of auto-labeling, surpassing manual annotation in some scenarios. The aforementioned methods are all designed for LiDAR point cloud, attributed to the ease of accumulating 3D spatial information compared to images. MV-Map [67] provides a region-centric fusion solution for high-definition maps with camera data as inputs while it only processes limited traffic elements on the bird’s eye view plane. To address the above challenges, We design the first hybrid propagation pipeline (OccFiner) for offboard semantic scene completion (SSC). OccFiner can process the SSC predictions from either cameras or LiDAR and provide high-quality occupancy information for autonomous vehicles, which holds significant importance in auto-labeling and data closure in pure visual perception.
3 Approach
3.1 Overview
@@
Figure 2: Overview of the proposed multi-to-multi local propagation network. It accepts multiple onboard predictions and relative coordinates as input. This network adeptly executes error compensation and facilitates the implicit local propagation, improving SSC quality across various distances.
3.2 Multi-to-Multi Local Propagation Network
Problem formulation: Let $X_{1}^{T}:=\{X_{1},X_{2},...,X_{T}\}$ be a onboard SSC inference sequence of height $H$, width $W$, depth $Z$, and frames length $T$. $C_{1}^{T}:=\{C_{1},C_{2},...,C_{T}\}$ denotes the corresponding frame-wise relative coordinates, which can be calculated by:
$\mathbf{c}_{i}=(T_{li}^{cam})^{\intercal}(T_{cam_{t}}^{world})^{\intercal}T_{cam_{i}}^{world}T_{li}^{cam}\mathbf{x}_{i},$ (1)
where each vertex $\mathbf{x}_{i}\in X_{i}$ is projectively associated to a relative vertex $\mathbf{c}_{i}\in C_{i}$. For each relative coordinate $C_{i}$, it represents the mapping relationship between the three-dimensional coordinates of the current frame $X_{i}$ and the coordinates of the pivot frame $X_{t}$. We formulate the occupancy refinement as a task that takes the onboard ($X_{1}^{T},C_{1}^{T}$) pairs as input and reconstruct the SSC labels $Y_{1}^{T}=\{Y_{1},Y_{2},...,Y_{T}\}$. Specifically, we propose to learn a mapping function from onboard inferences $X_{1}^{T}$ to the output $\hat{Y}_{1}^{T}:=\{\hat{Y}_{1},\hat{Y}_{2},...,\hat{Y}_{T}\}$, such that the conditional distribution of the real data $p(Y_{1}^{T}|X_{1}^{T})$ can be approximated by the one of generated data $p(\hat{Y}_{1}^{T}|X_{1}^{T})$.
We conceptualize the task as a “multi-to-multi” prediction problem, focusing on the simultaneous refinement of all input SSC frames in a single feed-forward process. The intuition is that regions obscured in the current viewpoint are likely to become visible in frames captured from a distance, particularly in scenarios involving large occlusions or when the vehicle is moving at a slower pace. In the context of multi-frame offboard SSC, it is more effective to address gaps in a target frame by leveraging content from the entire scene sequence, incorporating information from both neighboring and distant frames as conditional inputs. This approach relies on the Markov assumption [13] for simplification, which allows us to express the refinement process as:
$p(\hat{Y}_{1}^{T}|X_{1}^{T})=\prod_{t=1}^{T}p(\hat{Y}_{t-n}^{t+n}|X_{t-n}^{t+n},X_{t-d,s}^{t+d}),$ (2)
@@ 3.2 Multi-to-Multi Local Propagation Network
Network design: The overview of the proposed multi-to-multi local propagation network is shown in Fig. 2. As indicated in Eq. 2, OccFiner takes both neighboring local SSC frames $X_{t-n}^{t+n}$, distant reference SSC frames $X_{t-d,s}^{t+d}$, and relative coordinate $\{C_{t-n}^{t+n},C_{t-d,s}^{t+d}\}$ as conditions, to propagate across all input frames simultaneously. Specifically, OccFiner is comprised of four integral components, including a frame-level BEV feature encoder, a parallel BEV positional encoder, a series of spatial-temporal transformers, and a frame-level segmentation head. The frame-level BEV feature encoder is architecturally configured with multiple 2D convolutions along the horizontal and vertical dimensions (X, Y), thus converting the height dimension (Z) into a feature dimension. Mirroring this, the frame-level BEV positional encoder shares the same architectural blueprint with the feature encoder, such a design choice aids in better aligning the positional and feature information, as both are processed through a similar computational pathway. At the heart of OccFiner lie the DualFlow4D transformers, tasked with learning and applying joint spatial-temporal transformations. These transformers are specifically tailored to address uncertainties and occlusions within the deep encoding space.
@@
DualFlow4D transformer: To propagate high-fidelity features in each SSC frame, multi-layer DualFlow4D transformers are designed to search coherent contents from all the encoded features. Specifically, we propose to first search by a multi-head soft-patch-based focal attention module along BEV and temporal dimensions. We use $f_{1}^{T}=\{f_{1},f_{2},...,f_{T}\}$, where $f_{i}\in\mathbb{R}^{h\times w\times c}$ to denote the features encoded from the frame-level BEV encoder. Concurrently, $p_{l}\in\mathbb{R}^{T_{l}\times h\times w\times c}$ and $p_{r}\in\mathbb{R}^{T_{r}\times h\times w\times c}$ encapsulate the corresponding encoded local and reference relative coordinates. We first use soft split to embed them into overlapped patches $P\in\mathbb{R}^{(T_{l}+T_{r})\times W_{h}\times W_{w}\times c_{e}}$:
$P={\rm SS}(({f}_{l}\oplus f_{r})+({p}_{l}\oplus p_{r})),$ (3)
@@
where ${\rm SS}(\cdot)$ denotes the soft split operation, which softly splits each BEV feature into overlapped patches of size $W_{h}\times W_{w}\times T$ with stride $k$, and flattened to a one-dimensional token, which is similar to the image splitting strategy in T2T-ViT [76]. $T_{l}$ and $T_{r}$ are the time dimension of local frames and reference frames, and $W_{h}$ and $W_{w}$ are the spatial dimension of embedded tokens. $\oplus$ denotes the feature concatenation. The overlapped position aggregates a piece of information from different tokens, contributing to smoother semantic patch boundaries and enlarging its receptive field by fusing cues from neighboring patches.
Instead of the vanilla vision transformer [9], we use focal transformer [72] to search from both local and non-local neighbors to propagate high-fidelity BEV features. Concretely, $P$ is linearly projected to queries $Q$, keys $K$, and values $V$ for computing the 3D focal attention:
$Q,K,V=\mathcal{P}_{qkv}({\rm LN}(P)),$ (4)
$Z_{bev}={\rm MHFA}(Q,K,V)+P,$
where ${\rm LN}$ and ${\rm MHFA}$ denote the layer normalization and multi-head focal attention, respectively, $\mathcal{P}_{qkv}$ is the linear projection layer. We omit the time dimension for simplicity. The previous propagation is limited to the temporal BEV space. To ensure effective propagation in the height direction, we have additionally introduced attention along the z-axis, specifically focusing on the ‘pillar’, which refers to a voxel grid oriented along the height:
$Z_{z}:={\rm Attn}(Q_{z},K_{z},V_{z})={\rm Softmax}(\frac{Q_{z}K_{z}^{\top}V_{z}}{\sqrt{d}}),$ (5)
where $Q_{z},K_{z},V_{z}\in\mathbb{R}^{W_{z}\times T\times c_{e}}$ are respectively reshaped queries, keys, and values. Finally, the output of the entire propagated feature will be gathered:
$Z_{4d}=\mathcal{P}_{z}(Z_{z})+\mathcal{P}_{bev}(Z_{bev}),$ (6)
@@
where $\mathcal{P}_{z}$ and $\mathcal{P}_{bev}$ are linear projections. Note that the above formulas omit the head dimension for simplicity. Such a 4D dual-branch design promotes learning coherent spatial-temporal transformations by preserving both BEV information flow and the height dimension information flow along the pillar-shape voxels.
Figure 4: Overview of the proposed region-centric global propagation. Given an arbitrary number of input frames, we leverage multi-view geometry and consider sensor bias to augment long-term SSC consistency and accuracy.
3.3 Region-centric Global Propagation

### v4

# Offboard Occupancy Refinement with Hybrid Propagation for Autonomous Driving

@@ Abstract
Vision-based occupancy prediction, also known as 3D Semantic Scene Completion (SSC), presents a significant challenge in computer vision. Previous methods, confined to onboard processing, struggle with simultaneous geometric and semantic estimation, continuity across varying viewpoints, and single-view occlusion. Our paper introduces OccFiner, a novel offboard framework designed to enhance the accuracy of vision-based occupancy predictions. OccFiner operates in two hybrid phases: 1) a multi-to-multi local propagation network that implicitly aligns and processes multiple local frames for correcting onboard model errors and consistently enhancing occupancy accuracy across all distances. 2) the region-centric global propagation, focuses on refining labels using explicit multi-view geometry and integrating sensor bias, particularly for increasing the accuracy of distant occupied voxels. Extensive experiments demonstrate that OccFiner improves both geometric and semantic accuracy across various types of coarse occupancy, setting a new state-of-the-art performance on the SemanticKITTI dataset. Notably, OccFiner significantly boosts the performance of vision-based SSC models, achieving accuracy levels competitive with established LiDAR-based onboard SSC methods. Furthermore, OccFiner is the first to achieve automatic annotation of SSC in a purely vision-based approach. Quantitative experiments prove that OccFiner successfully facilitates occupancy data loop-closure in autonomous driving. Additionally, we quantitatively and qualitatively validate the superiority of the offboard approach on city-level SSC static maps. The source code will be made publicly available at https://github.com/MasterHow/OccFiner.
Index Terms:
Semantic Scene Completion, 3D Occupancy Prediction, Offboard Perception, Data Loop-Closure.
I Introduction
@@
Vision-based Occupancy Prediction, often referred to as Semantic Scene Completion (SSC), aims to accurately reconstruct the geometry and semantics of a 3D scene from image captures [1], providing critical location and semantic information for safe navigation in autonomous driving. Compared to LiDAR solutions, camera-based systems offer the benefits of being lightweight, cost-effective, and easy to deploy and maintain.
@@
However, vision-based SSC has historically lagged in accuracy compared to LiDAR due to two main limitations: 1) The depth measurement of monocular cameras lacks the precision of LiDAR systems, requiring onboard vision-based SSC methods to struggle with both geometry and semantics. 2) Onboard models are limited to processing a restricted number of local frames, impacting accuracy and multi-view continuity. This issue manifests as unreliable estimations and discontinuities across viewpoints in geometry and semantics. As illustrated in Fig. 1, most SSC efforts are focused on onboard settings [2, 1, 3, 4, 5, 6, 7, 8], i.e., data calculation and feedback occur on the vehicle itself. However, integrating data into offboard systems is critical for identifying and rectifying potential model deficiencies, achieving data loop-closure, and enabling system self-evolution, which are essential for a wide range of intelligent transportation applications [9]. This approach solves the distribution shift between training and deployment and allows training to be scaled both safely and cost-effectively [10]. While offboard systems offer refined evaluation of onboard model performance, essential for the data loop-closure in autonomous driving [10, 9, 11], and well-studied in 3D detections [12, 13, 14, 15], their potential in SSC remains under-explored.
Given such limitations of inferior performance and onboard SSC settings, we propose the first offboard SSC framework, OccFiner, a novel approach that involves additional computation and refinement of on-vehicle predictions at a data center but greatly boosts the prediction reliability. The objective is to significantly enhance the reliability of vision-based SSC. By doing so, the constructed 3D SSC map can be reliably used repeatedly after a one-time low-cost acquisition with only cameras. This approach enhances auto-labeling by leveraging aggregated and diverse data from multiple sources, significantly reducing manual annotation efforts and improving labeling accuracy for large-scale 3D scene understanding [10, 9, 11] in autonomous driving.
@@
Specifically, we consider two primary error sources in addressing the inaccuracies of vision-based SSC: the prediction bias error introduced by the onboard model, and the measurement bias from the camera capture process. OccFiner decouples these two heterogeneous error sources and uses a two-stage hybrid process to process them separately: 1) Multi-to-Multi Local Propagation. In the initial stage, OccFiner addresses prediction bias errors from the onboard model. We implement an error compensation model that processes inputs from off-the-shelf SSC models [1, 4, 2]. This model uniquely encodes relative spatial coordinates within a local window, facilitating the propagation of geometric and semantic cues by a newly proposed DualFlow4D transformer. This approach effectively enhances the accuracy of occupancy predictions at varying distances. 2) Region-centric Global Propagation. The second stage focuses on measurement errors from camera captures. Here, OccFiner employs explicit multi-view geometry to register and aggregate data across the entire scene, incorporating sensor bias weighting in the process. Compared with the simple accumulation of multi-frame averages, our region-centric voting strategy yields a more accurate SSC map. By integrating local implicit error compensation and non-local explicit information propagation from various viewpoints, OccFiner effectively addresses onboard bias, multi-view consistency, and single-view occlusions, thereby naturally enhancing the reliability and robustness of scene understanding.
TABLE I: Glossary of Terms and Abbreviations.
Abbreviation Full Form
SSC Semantic Scene Completion
BEV Bird’s-Eye View
V2X Vehicle-to-Everything
SS Soft Split Operation
LN Layer Normalization
Q Queries
K Keys
V Values
MHFA Multi-Head Focal Attention
Attn Scaled Dot-product Attention
DualFlow4D Dual-branch 4D Feature Flow Module
IoU Intersection over Union
mIoU Mean Intersection over Union
fov Field of View
@@ V Values
To the best of our knowledge, we are the first to introduce offboard refinement for SSC. Our method, termed OccFiner, enables seamless integration across various SSC models and datasets as a plug-and-play solution. For 3D semantic scene completion on the SemanticKITTI dataset [16], OccFiner has demonstrated remarkable performance, surpassing VoxFormer by $3.67\%$ in mIoU, which makes a $26.99\%$ relative improvement and ranks first on the test leaderboard among all vision-based methods. Surprisingly, OccFiner also works well for LiDAR-based SSC [17] with minor changes, leading to a new state-of-the-art of $37.82\%$ in mIoU, achieving a $9.5\%$ improvement compared with the best benchmark result.
To enhance readability, Tab. I presents all abbreviations and their full forms used in this paper.
@@
  • We introduce OccFiner, the first dedicated offboard framework designed to enhance vision-based Semantic Scene Completion (SSC) by leveraging multi-view consistency and addressing onboard model limitations.
  • We propose a novel hybrid propagation strategy within OccFiner, uniquely combining multi-frame local implicit refinement (via DualFlow4D transformer) to correct prediction bias and global explicit aggregation (via region-centric voting with sensor bias) to handle measurement bias and ensure long-range consistency.
  • OccFiner achieves new state-of-the-art performance on the public SemanticKITTI benchmark for both camera-based and LiDAR-based SSC. Notably, our refined camera-based SSC is comparable with established LiDAR-based onboard SSC methods (e.g. SSCNet-full [18]), demonstrating substantial relative mIoU improvements (e.g., $+24.59\%$ refinement for VoxFormer on the validation set).
  • We successfully demonstrate the first application of purely vision-based automatic annotation for SSC data loop-closure, validating OccFiner’s potential for scalable data generation and system self-evolution.
  • We validate OccFiner’s effectiveness on city-level mapping, proving the benefits of the offboard approach for generating large-scale, consistent 3D semantic maps.
II Related Work
II-A 3D Semantic Scene Completion
@@ II Related Work
II-B Offboard 3D Perception
Perception in autonomous driving scenarios necessitates extensive data annotation and training [69, 70, 71]. Beyond the real-time ‘onboard’ processing performed directly on the vehicle, ‘offboard’ perception involves processing collected sensor data after the driving mission, typically within a data center or cloud environment. This paradigm leverages significantly larger computational resources and allows for the aggregation of data from multiple vehicles or extensive driving logs. Key goals of offboard systems include identifying and rectifying potential onboard model deficiencies, achieving data loop-closure [9, 10, 11], enabling large-scale automatic annotation, and facilitating system self-evolution [9], which are crucial for scaling autonomous driving capabilities safely and cost-effectively [10]. Since LiDAR can provide accurate range information in 3D space, researchers explore its sequence learning in semantic segmentation [72, 73], object detection [74, 75], and tracking [76] to obtain better performance. The sequential characteristics of LiDAR scans are further studied in offboard 3D object detection [12, 13, 14, 15] to reduce annotation costs. 3DAL [12] formulates the problem of offboard 3D object detection and proposes an object-centric auto-labeling detector. CTRL [14] adheres to the principle of track-centric and elevates the accuracy of auto-labeling, surpassing manual annotation in some scenarios. While these aforementioned methods demonstrated the value of the offboard approach, they were primarily designed for LiDAR point clouds and focused on object detection or tracking. The challenge of applying offboard refinement to dense, vision-based Semantic Scene Completion (SSC), which involves reconstructing both geometry and semantics voxel-wise, remained under-explored. Vision-based approaches like MV-Map [77] provided region-centric fusion for high-definition maps but only processed limited traffic elements on the bird’s-eye view plane. To address these gaps and harness the benefits of offboard processing for dense scene understanding from cameras, we design the first hybrid propagation pipeline (OccFiner) specifically for offboard SSC. OccFiner can process the SSC predictions from either cameras or LiDAR and provide high-quality occupancy information for autonomous vehicles, which holds significant importance in auto-labeling and data closure in pure visual perception.
III Approach
III-A Overview
Figure 2: Overview of the proposed multi-to-multi local propagation network (OccFiner Stage 1). It accepts multiple onboard predictions and relative coordinates as input (multi-input) and simultaneously generates refined predictions for multiple frames (multi-output). This network adeptly executes error compensation and facilitates implicit local propagation, improving SSC quality and temporal consistency across various distances.
@@ III-A Overview
III-B Multi-to-Multi Local Propagation Network
Problem Formulation: Let $X^{T}:=\{X_{1},X_{2},...,X_{T}\}$ be an onboard SSC inference sequence of height $H$, width $W$, depth $Z$, and $T$ frames in length. $C^{T}:=\{C_{1},C_{2},...,C_{T}\}$ denotes the corresponding frame-wise relative coordinates, which can be calculated by:
$\mathbf{c}_{i}=(T_{li}^{cam})^{-1}(T_{cam_{t}}^{world})^{-1}T_{cam_{i}}^{world}T_{li}^{cam}\mathbf{x}_{i},$ (1)
where each vertex $\mathbf{x}_{i}\in X_{i}$ is projectively associated with a relative vertex $\mathbf{c}_{i}\in C_{i}$. For each relative coordinate $C_{i}$, it represents the mapping relationship between the three-dimensional coordinates of the current frame $X_{i}$ and the coordinates of the pivot frame $X_{t}$. We formulate the occupancy refinement as a task that takes the onboard ($X^{T},C^{T}$) pairs as input and reconstruct the SSC labels $Y^{T}=\{Y_{1},Y_{2},...,Y_{T}\}$. Specifically, we propose to learn a mapping function from onboard inferences $X^{T}$ to the output $\hat{Y}^{T}:=\{\hat{Y}_{1},\hat{Y}_{2},...,\hat{Y}_{T}\}$, such that the conditional distribution of the real data $p(Y^{T}|X^{T})$ can be approximated by the distribution of generated data $p(\hat{Y}^{T}|X^{T})$.
We conceptualize the task as a ‘multi-to-multi’ prediction problem: the network takes multiple input frames ($X^{T}$) and simultaneously refines multiple output frames ($\hat{Y}^{T}$) within that sequence in a single feed-forward pass. This approach is chosen specifically to leverage temporal context and enforce consistency across the refined sequence, unlike methods that process frames individually or refine only a single central frame. The intuition is that regions obscured in the current viewpoint are likely to become visible in frames captured from a distance, particularly in scenarios involving large occlusions or when the vehicle is moving at a slower pace. In the context of multi-frame offboard SSC, it is more effective to address gaps in a target frame by leveraging content from the entire scene sequence, incorporating information from both neighboring and distant frames as conditional inputs. This approach relies on the Markov assumption [78] for simplification, which allows us to express the refinement process as:
$p(\hat{Y}^{T}|X^{T})=\prod_{t=1}^{T}p(\hat{Y}_{t-n}^{t+n}|X_{t-n}^{t+n},X_{t-d,s}^{t+d}),$ (2)
@@ III-B Multi-to-Multi Local Propagation Network
Network Design: The overview of the proposed multi-to-multi local propagation network is shown in Fig. 2. As indicated in Eq. 2, OccFiner takes both neighboring local SSC frames $X_{t-n}^{t+n}$, distant reference SSC frames $X_{t-d,s}^{t+d}$, and relative coordinate $\{C_{t-n}^{t+n},C_{t-d,s}^{t+d}\}$ as conditions, to propagate across all input frames simultaneously. Specifically, OccFiner comprises four integral components, including a frame-level Bird’s-Eye View (BEV) feature encoder, a parallel BEV positional encoder, a series of spatial-temporal transformers, and a frame-level segmentation head. The frame-level BEV feature encoder is architecturally configured with multiple 2D convolutions along the horizontal and vertical dimensions (X, Y), thus converting the height dimension (Z) into a feature dimension. Mirroring this, the frame-level BEV positional encoder shares the same architectural blueprint with the feature encoder; such a design choice aids in better aligning the positional and feature information, as both are processed through a similar computational pathway. At the heart of OccFiner lie the DualFlow4D transformers, tasked with learning and applying joint spatial-temporal transformations. These transformers are specifically tailored to address uncertainties and occlusions within the deep encoding space.
@@
DualFlow4D Transformer: To propagate high-fidelity features in each SSC frame, multi-layer DualFlow4D transformers are designed to search for coherent from all the encoded features. Specifically, we propose to first search by a multi-head soft-patch-based focal attention module along BEV and temporal dimensions. We use $f^{T}=\{f_{1},f_{2},...,f_{T}\}$ to denote the encoded BEV features, where $f_{i}\in\mathbb{R}^{h\times w\times c}$, and $T$ is the total of local and reference frames. Concurrently, $p_{l}\in\mathbb{R}^{T_{l}\times h\times w\times c}$ and $p_{r}\in\mathbb{R}^{T_{r}\times h\times w\times c}$ encapsulate the corresponding encoded local and reference relative coordinates, where $T_{l}$ and $T_{r}$ are the time dimension of local frames and reference frames. Let ${\rm SS}(\cdot)$ denote the soft split operation, which reshapes features into overlapping patches. Let $\oplus$ denote the feature concatenation along the channel dimension. We first embed the features and coordinates into overlapped patches $P\in\mathbb{R}^{(T_{l}+T_{r})\times W_{h}\times W_{w}\times C_{e}}$ as follows:
$P={\rm SS}(({f}_{l}\oplus f_{r})+({p}_{l}\oplus p_{r})).$ (3)
@@
Here, $f_{l}$ and $f_{r}$ represent the features from local and reference frames, respectively. $W_{h}$ and $W_{w}$ are the spatial dimensions of the embedded tokens, and $C_{e}$ is the embedding channel dimension. The overlapped position aggregates a piece of information from different tokens, contributing to smoother semantic patch boundaries and enlarging its receptive field by fusing cues from neighboring patches.
Instead of the vanilla vision transformer [79], we use the focal transformer mechanism [80] to efficiently aggregate information from local and non-local neighboring tokens and propagate high-fidelity BEV features. Let $P$ be the sequence of embedded patches from Eq. (3). We first apply Layer Normalization (${\rm LN}$) to $P$. Then, a linear projection layer, denoted as $\mathcal{P}_{qkv}$, maps the normalized tokens to Queries ($Q$), Keys ($K$), and Values ($V$):
$Q,K,V=\mathcal{P}_{qkv}({\rm LN}(P)).$ (4)
These $Q$, $K$, and $V$ are then processed by the Multi-Head Focal Attention module (${\rm MHFA}$). The resulting aggregated feature for the BEV space, $Z_{bev}$, is obtained by adding the original input $P$ (residual connection):
$Z_{bev}={\rm MHFA}(Q,K,V)+P.$ (5)
For simplicity in notation, we omit the explicit time dimension here. The propagation described by Eq. (5) operates primarily within the temporal BEV space. To ensure effective feature propagation along the height dimension (Z-axis), we introduce an additional attention mechanism. This focuses on ‘pillars,’ where each pillar represents a voxel stack along the height axis at a specific BEV location ($x,y$). We reshape the features to obtain pillar-specific Queries ($Q_{z}$), Keys ($K_{z}$), and Values ($V_{z}$), typically with dimensions like $\mathbb{R}^{W_{z}\times T\times c_{e}}$. Let $d$ be the dimension of the key vectors. We compute the attention output along the pillars, $Z_{z}$, using standard scaled dot-product attention (${\rm Attn}$):
$Z_{z}:={\rm Attn}(Q_{z},K_{z},V_{z})={\rm Softmax}(\frac{Q_{z}K_{z}^{\top}V_{z}}{\sqrt{d}}).$ (6)
Having computed the BEV-focused features $Z_{bev}$ and the pillar-focused features $Z_{z}$, we combine them to form the final output of the DualFlow4D block. Let $\mathcal{P}_{z}$ and $\mathcal{P}_{bev}$ be linear projection layers specific to the pillar and BEV branches, respectively. The final 4D propagated feature, $Z_{4d}$, is obtained by projecting each branch’s output and summing them:
$Z_{4d}=\mathcal{P}_{z}(Z_{z})+\mathcal{P}_{bev}(Z_{bev}).$ (7)
@@
Note that we omit the multi-head dimension in the notation for simplicity. This 4D dual-branch design promotes learning coherent spatial-temporal transformations by explicitly processing both BEV information flow ($Z_{bev}$) and height dimension information flow ($Z_{z}$) along the pillar-shape voxels.
Figure 4: Overview of the OccFiner two-stage pipeline. Stage 1 performs Multi-to-Multi Local Propagation to refine onboard SSC predictions ($X^{T}$) into $\hat{Y}^{T}$. Stage 2 then performs Region-centric Global Propagation on $\hat{Y}^{T}$, leveraging multi-view geometry and sensor-aware voting to generate the final consistent and accurate offboard SSC map, which can be further applied to auto labeling or data loop-closure.
III-C Region-centric Global Propagation